# From Cursed to Competitive: Closing the ZO–FO Gap via Input-to-State Stability

### 28 Apr 2026 v1

@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is continuously differentiable, integrable, and bounded below. Such problems arise routinely in machine learning, signal processing, and control. While previously, in the literature, it was commonly understood that convergence complexity bounds for zeroth-order (ZO) methods have extra dimension dependencies compared to their first-order (FO) counterparts, in this work, we analyse conditions under which it is guaranteed that ZO methods, in expectation, have no extra dimension dependencies in their convergence rate compared to their FO counterparts. Here, we study iterative algorithms and their convergence properties from the dynamical system perspective and analyse them by studying input-to-state stability (ISS) properties.
Motivation: In many modern applications, gradient information is either unavailable, prohibitively expensive to compute, or simply does not exist in closed form. Examples include black-box adversarial attacks (Chen et al., 2017; Ye et al., 2018; Madry et al., 2017), reinforcement learning with opaque simulators (Salimans et al., 2017; Choromanski et al., 2018), hyperparameter tuning (Snoek et al., 2012), simulation-based engineering design (Conn et al., 2009), and memory-efficient fine-tuning of large language models (Malladi et al., 2023). In these settings, one must resort to ZO methods that construct gradient surrogates from function evaluations alone. While attractive for their simplicity and broad applicability, ZO methods have traditionally been seen as having fundamentally weaker convergence guarantees than their FO counterparts, independent of the parameter selection.
@@
 (ii) Leveraging the ISS framework, we prove that ZO methods inherit the same transient convergence rate as their FO counterparts, i.e., the same exponential/asymptotic decay, and converge to a neighbourhood of the FO fixed point whose radius depends on the perturbation bound but carries no additional dependence on the number of iterations (Theorems 1 and 2). This provides a fundamentally different explanation of ZO convergence behaviour. The dimension dependence does not appear in the iteration complexity but rather in the size of the convergence neighbourhood. Moreover, the framework serves as a general-purpose tool. Whenever the FO algorithm is known to converge, the ISS machinery automatically yields convergence of its ZO counterpart to a neighbourhood of the FO fixed point, without requiring a separate, algorithm-specific proof.
@@
 (iv) We validate the theoretical findings through numerical experiments, confirming that ZO trajectories closely track their FO counterparts with no extra dimension dependence, and demonstrating how $\mu$, $h$, and $t$ control this proximity.
@@
Notation: For $x,y\in\mathbb{R}^{d}$, we write $\langle x,y\rangle=x^{\top}y$, and let $\|\cdot\|$ denote the Euclidean norm of a vector and the corresponding spectral norm of a matrix. For $\delta>0$ and $z^{\ast}\in\mathbb{R}^{d}$, we define $\mathcal{B}_{\delta}(z^{\ast}):=\{z\in\mathbb{R}^{d}\mid\|z-z^{\ast}\|\leq\delta\}$. For $f:\mathbb{R}^{n}\to\mathbb{R}$, we denote the gradient and Hessian by $\nabla f$ and $\nabla^{2}f$, respectively. The identity matrix of proper size is denoted by $\mathrm{I}$, and $\mathrm{Id}$ denotes the identity function $\mathrm{Id}(s)=s$. A continuous function $\alpha:[0,a)\to[0,\infty)$ (with $a>0$ or $a=\infty$) is of class $\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$. A function $\alpha$ is of class $\mathcal{K}_{\infty}$ if $\alpha\in\mathcal{K}$ and $\alpha(s)\to\infty$ as $s\to\infty$. A continuous function $\beta:[0,\infty)\times[0,\infty)\to[0,\infty)$ is of class $\mathcal{KL}$ if, for each fixed $t\geq 0$, the map $s\mapsto\beta(s,t)$ belongs to $\mathcal{K}$, and for each fixed $s\geq 0$, the map $t\mapsto\beta(s,t)$ is decreasing with $\beta(s,t)\to 0$ as $t\to\infty$.
@@ 2.1 Dynamical systems and perturbations
with state $z\in\mathbb{R}^{d}$ and update map $w:\mathbb{R}^{d}\to\mathbb{R}^{d}$. A point $z^{e}\in\mathbb{R}^{d}$ is an equilibrium (fixed point) of $w$ if $w(z^{e})=z^{e}$. To analyse stability of equilibria, we use standard notions from discrete-time systems theory; see, e.g., (Kellett and Braun, 2023, Ch. 2, Ch. 5).
@@
An equilibrium $z^{e}\in\mathbb{R}^{d}$ of (2) is stable if there exist $\alpha\in\mathcal{K}_{\infty}$ and $\delta>0$ such that $\|z_{k}-z_{e}\|\leq\alpha(\|z_{0}-z_{e}\|)$ for all $k\in\mathbb{N}$ and $z_{0}\in\mathcal{B}_{\delta}(z^{e})$. If $z^{e}$ is not stable, it is called unstable. The equilibrium is asymptotically stable if there exist $\beta\in\mathcal{KL}$ and $\delta>0$ such that
@@
If $\beta$ in (3) has the form $\beta(r,k)=Mr\gamma^{k}$ for $M>0$ and $\gamma\in(0,1)$, then $z^{e}$ is exponentially stable.
@@
The system (4) with $z_{k},q_{k}\in\mathbb{R}^{d}$ is (locally) input-to-state stable (ISS) if there exist $\delta>0$, $\beta\in\mathcal{KL}$, and $\gamma\in\mathcal{K}$ such that, for all sequences $\{q_{k}\}_{k\geq 0}$,
@@ 2.2 Gaussian smoothing
This is a standard assumption; see, e.g., Maass et al. (2021); Liu et al. (2020); Farzin et al. (2025a). The variance can be reduced by averaging $t$ independent oracle evaluations per iteration, yielding $E_{u}\!\left[\|g_{\mu}(x_{k})-\nabla f_{\mu}(x_{k})\|^{2}\right]\leq\sigma^{2}/t$ while preserving unbiasedness (Balasubramanian and Ghadimi, 2022). The details of relaxation of Assumption 1 to a more relaxed notion are given in Remark 4.
@@ 3.1 Gradient descent and strong convexity
Lemma 2 shows that the averaged ZO-GD iterates satisfy a perturbed version of the GD dynamics, where the perturbations are bounded and their magnitude can be manipulated through appropriate parameter selection. It is known that, for a smooth strongly convex objective, GD converges to the minimiser exponentially fast Nesterov and others (2018). In Nesterov and Spokoiny (2017), ZO-GD was analysed under the same setting, and it was shown that the iteration complexity carries an extra dimension-dependent factor for any choice of parameters. The following theorem shows that, instead, the dimension affects only the size of the convergence neighbourhood, and there are parameter choices such that ZO-GD closely follows GD.
@@ 3.2 Heavy ball method and Nesterov’s accelerated gradient under strong convexity
$\sup_{k}E[\|q_{k}\|]\leq\frac{h_{1}\mu L_{1}(n+3)^{3/2}}{2}+\frac{h_{1}^{2}L_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{HB}}}}\,.$ (26)
@@
$\sup_{k}E[\|q_{k}\|]\leq\frac{h_{1}\mu L_{1}(n+3)^{3/2}}{2}+\frac{(1+2h_{2})h_{1}^{2}L_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{NAG}}}}\,.$ (28)
@@
Both lemmas show that the averaged ZO iterates satisfy a perturbed version of their FO dynamics with bounded, controllable perturbations. The key structural difference is that, for NAG, the strong convexity contraction acts directly on the lookahead error, eliminating the momentum cross term that appears in the HB analysis. For HB, it is well known that the method locally converges exponentially fast Polyak (1964); for NAG, global convergence with an accelerated rate is established in Nesterov and others (2018). The following theorems provide a unified perspective that for both methods, the dimension dependence does not appear in the iteration complexity, instead, the convergence is to a neighbourhood of FO dynamics, where the size of the neighbourhood can be made arbitrarily small.
@@
Consider (22) and (23) with $P=T$. Let $f$ be $\beta$-strongly convex with $L_{1}$-Lipschitz gradients, let the hypotheses of Lemma 3 be satisfied, and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of HB. Then there exist positive scalars $\bar{\mu}$ and $\bar{q}$, and $\bar{t}\in\mathbb{N}$ such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}E[\|q_{k}\|]\leq\bar{q}$ via the bound in Lemma 3, ZO-HB has the same exponential decay rate as HB and converges to $\mathcal{B}_{r}(z^{e})$, and there exists a class-$\mathcal{K}$ function $\alpha$ such that
@@
Consider (22) and (23) with $P=TA$. Let $f$ be $\beta$-strongly convex with $L_{1}$-Lipschitz gradients, let the hypotheses of Lemma 4 be satisfied, and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of NAG. Then there exist $\bar{\mu},\bar{q}>0$ and $\bar{t}\in\mathbb{N}$ such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}E[\|q_{k}\|]\leq\bar{q}$ via the bound in Lemma 4, ZO-NAG has the same decay rate as NAG and converges to $\mathcal{B}_{r}(z^{e})$, and there exists a class-$\mathcal{K}$ function $\alpha$ such that
@@ 3.3 $L_{2}$ regularisation
In the preceding sections, we showed that strong convexity is sufficient to guarantee that the averaged ZO dynamics can be written as a bounded perturbation of the FO dynamics. In this section, we show that adding an $L_{2}$ regularisation term to the problem (1) provides sufficient contraction to recover the same perturbation framework, even when $f$ itself is not strongly convex.
@@
Consider (32), (34), and (35). Let $f$ be a possibly non-strongly-convex function with Lipschitz gradients with constant $L_{1}>0$, and let $h\in\bigl(0,\frac{\sqrt{2-c^{2}}-c}{4L_{1}}\bigr)$. Then
$\sup_{k}\|q_{k}\|\leq h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-8L_{1}^{2}h^{2})}}.$ (36)
Remark 1.
Since $f$ has $L_{1}$-Lipschitz gradients, regularising with $\lambda>\frac{L_{1}}{2}$ makes the overall objective strongly convex, and one could directly apply the results of Section 3.1. With the parametrisation $\lambda=\frac{1-c}{2h}$, where $c\in[0,1)$ and $h\in\bigl(0,\frac{\sqrt{2-c^{2}}-c}{4L_{1}}\bigr)$, we have $\lambda>\frac{2L_{1}(1-c)}{\sqrt{2-c^{2}}-c}\geq\sqrt{2}L_{1}(1-c)$. Consequently, for $c\leq 1-\frac{1}{2\sqrt{2}}$ the regularised objective is strongly convex, whereas for $c>1-\frac{1}{2\sqrt{2}}$ it need not be. This reveals a trade-off governed by $c.$ A small value of $c$ yields strong convexity, a larger admissible range for $h$, and more control over the magnitude of the perturbations, but it also introduces a larger shift of the minimiser from that of the original problem. Conversely, a large value of $c$ preserves proximity to the original minimiser and still provides sufficient contraction for the perturbation to remain bounded, at the expense of a smaller admissible $h$. Thus, if fidelity to the original minimiser is the priority, one should choose $c$ large and control the perturbation through $t$ and $h$; if stability properties are more important, one can choose $c$ smaller and benefit from smaller bounds on the perturbation.
@@
Consider (32) and (34). Let $f$ be a possibly non-strongly-convex function with Lipschitz gradients with constant $L_{1}>0$, and let $h\in\bigl(0,\frac{\sqrt{2-c^{2}}-c}{4L_{1}}\bigr),$ and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of GD. Suppose that GD applied to the regularised problem (31) converges to a fixed point $z^{e}$. Then there exist constants $\bar{\mu},\bar{q}>0$ and $\bar{t}\in\mathbb{N}$ (depending on $h$, $c$, $L_{1}$, and the Lyapunov function of the GD dynamics) such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}\|q_{k}\|\leq\bar{q}$ via the bound in Lemma 5, ZO-GD has the same decay rate as GD and converges to $\mathcal{B}_{r}(z^{e})$ and there exists a class-$\mathcal{K}$ function $\alpha$ such that
$r=\alpha\left(h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-8L_{1}^{2}h^{2})}}\right).$ (37)
In the next section, we illustrate the theoretical findings through numerical examples.

### 01 May 2026 v2

@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is continuously differentiable, and bounded below. Such problems arise routinely in machine learning, signal processing, and control. While previously, in the literature, it was commonly understood that convergence complexity bounds for zeroth-order (ZO) methods have extra dimension dependencies compared to their first-order (FO) counterparts, in this work, we analyse conditions under which it is guaranteed that ZO methods, in expectation, have no extra dimension dependencies in their convergence rate compared to their FO counterparts. Here, we study iterative algorithms and their convergence properties from the dynamical system perspective and analyse them by studying input-to-state stability (ISS) properties.
Motivation: In many modern applications, gradient information is either unavailable, prohibitively expensive to compute, or simply does not exist in closed form. Examples include black-box adversarial attacks (Chen et al., 2017; Ye et al., 2018; Madry et al., 2018), reinforcement learning with opaque simulators (Salimans et al., 2017; Choromanski et al., 2018), hyperparameter tuning (Snoek et al., 2012), simulation-based engineering design (Conn et al., 2009), and memory-efficient fine-tuning of large language models (Malladi et al., 2023). In these settings, one must resort to ZO methods that construct gradient surrogates from function evaluations alone. While attractive for their simplicity and broad applicability, ZO methods have traditionally been seen as having fundamentally weaker convergence guarantees than their FO counterparts, independent of the parameter selection.
@@
 (ii) Leveraging the ISS framework, we prove that ZO methods inherit the same transient convergence rate as their FO counterparts, i.e., the same exponential/asymptotic decay, and converge to a neighbourhood of the FO fixed point whose radius depends on the perturbation bound but carries no additional dependence on the number of iterations (Theorems 1 and 2). This provides a fundamentally different explanation of ZO convergence behaviour. The dimension dependence does not appear in the iteration complexity but rather in the size of the convergence neighbourhood. Moreover, the framework serves as a general-purpose tool. Whenever the FO algorithm is known to converge, the ISS machinery automatically yields convergence of its ZO counterpart to a neighbourhood of the FO fixed point, without requiring a separate, algorithm-specific proof. Our results concern iteration complexity. The per-iteration cost of each ZO step involves $t$ function evaluations. It may depend on problem parameters, including the variance, to achieve a given neighbourhood radius.
@@
 (iv) We validate the theoretical findings through numerical experiments, confirming that ZO trajectories closely track their FO counterparts and demonstrating how $\mu$, $h$, and $t$ control the proximity.
@@
Notation: For $x,y\in\mathbb{R}^{n}$, we write $\langle x,y\rangle=x^{\top}y$, and let $\|\cdot\|$ denote the Euclidean norm of a vector and the corresponding spectral norm of a matrix. For $\delta>0$ and $z^{\ast}\in\mathbb{R}^{n}$, we define $\mathcal{B}_{\delta}(z^{\ast}):=\{z\in\mathbb{R}^{n}\mid\|z-z^{\ast}\|\leq\delta\}$. For $f:\mathbb{R}^{n}\to\mathbb{R}$, we denote the gradient and Hessian by $\nabla f$ and $\nabla^{2}f$, respectively. The identity matrix of proper size is denoted by $\mathrm{I}$, and $\mathrm{Id}$ denotes the identity function $\mathrm{Id}(s)=s$. A continuous function $\alpha:[0,a)\to[0,\infty)$ (with $a>0$ or $a=\infty$) is of class $\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$. A function $\alpha$ is of class $\mathcal{K}_{\infty}$ if $\alpha\in\mathcal{K}$ and $\alpha(s)\to\infty$ as $s\to\infty$. A continuous function $\beta:[0,\infty)\times[0,\infty)\to[0,\infty)$ is of class $\mathcal{KL}$ if, for each fixed $t\geq 0$, the map $s\mapsto\beta(s,t)$ belongs to $\mathcal{K}$, and for each fixed $s\geq 0$, the map $t\mapsto\beta(s,t)$ is decreasing with $\beta(s,t)\to 0$ as $t\to\infty$.
@@ 2.1 Dynamical systems and perturbations
with state $z\in\mathbb{R}^{n}$ and update map $w:\mathbb{R}^{n}\to\mathbb{R}^{n}$. A point $z^{e}\in\mathbb{R}^{n}$ is an equilibrium (fixed point) of $w$ if $w(z^{e})=z^{e}$. To analyse stability of equilibria, we use standard notions from discrete-time systems theory; see, e.g., (Kellett and Braun, 2023, Ch. 2, Ch. 5).
@@
An equilibrium $z^{e}\in\mathbb{R}^{n}$ of (2) is stable if there exist $\alpha\in\mathcal{K}_{\infty}$ and $\delta>0$ such that $\|z_{k}-z_{e}\|\leq\alpha(\|z_{0}-z_{e}\|)$ for all $k\in\mathbb{N}$ and $z_{0}\in\mathcal{B}_{\delta}(z^{e})$. If $z^{e}$ is not stable, it is called unstable. The equilibrium is asymptotically stable if there exist $\beta\in\mathcal{KL}$ and $\delta>0$ such that
@@
If $\beta$ in (3) has the form $\beta(r,k)=Mr\nu^{k}$ for $M>0$ and $\nu\in(0,1)$, then $z^{e}$ is exponentially stable.
@@
The system (4) with $z_{k},q_{k}\in\mathbb{R}^{n}$ is (locally) input-to-state stable (ISS) if there exist $\delta>0$, $\beta\in\mathcal{KL}$, and $\gamma\in\mathcal{K}$ such that, for all sequences $\{q_{k}\}_{k\geq 0}$,
@@ 2.2 Gaussian smoothing
This is a standard assumption; see, e.g., Maass et al. (2021); Liu et al. (2020); Farzin et al. (2025a). The variance can be reduced by averaging $t$ independent oracle evaluations per iteration, yielding $E_{u}\!\left[\|g_{\mu}(x_{k})-\nabla f_{\mu}(x_{k})\|^{2}\right]\leq\sigma^{2}/t$ while preserving unbiasedness (Balasubramanian and Ghadimi, 2022). We note that since our convergence results are local (Theorems 1 and 2 hold on a ball $B_{\bar{r}}(z^{e})$), Assumption 1 needs only to hold on this compact region. A relaxation to state-dependent variance is analysed in Remark 3.
@@ 3.1 Gradient descent and strong convexity
Lemma 2 shows that the averaged ZO-GD iterates satisfy a perturbed version of the GD dynamics, where the perturbations are bounded and their magnitude can be manipulated through appropriate parameter selection. It is known that, for a smooth strongly convex objective, GD converges to the minimiser exponentially fast Nesterov (2018). In Nesterov and Spokoiny (2017), ZO-GD was analysed under the same setting, and it was shown that the iteration complexity carries an extra dimension-dependent factor for any choice of parameters. The following theorem shows that, instead, the dimension affects only the size of the convergence neighbourhood, and there are parameter choices such that ZO-GD closely follows GD.
@@ 3.2 Heavy ball method and Nesterov’s accelerated gradient under strong convexity
$\sup_{k}\|q_{k}\|\leq\frac{h_{1}\mu L_{1}(n+3)^{3/2}}{2}+\frac{h_{1}^{2}L_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{HB}}}}\,.$ (26)
@@
$\sup_{k}\|q_{k}\|\leq\frac{h_{1}\mu L_{1}(n+3)^{3/2}}{2}+\frac{(1+2h_{2})h_{1}^{2}L_{1}\sigma}{\sqrt{t}}\sqrt{\frac{2}{1-\rho_{\mathrm{NAG}}}}\,.$ (28)
@@
Both lemmas show that the averaged ZO iterates satisfy a perturbed version of their FO dynamics with bounded, controllable perturbations. The key structural difference is that, for NAG, the strong convexity contraction acts directly on the lookahead error, eliminating the momentum cross term that appears in the HB analysis. For HB, it is well known that the method locally converges exponentially fast Polyak (1964); for NAG, global convergence with an accelerated rate is established in Nesterov (2018). The following theorems provide a unified perspective that for both methods, the dimension dependence does not appear in the iteration complexity, instead, the convergence is to a neighbourhood of FO dynamics, where the size of the neighbourhood can be made arbitrarily small.
@@
Consider (22) and (23) with $P=T$. Let $f$ be $\beta$-strongly convex with $L_{1}$-Lipschitz gradients, let the hypotheses of Lemma 3 be satisfied, and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of HB. Then there exist positive scalars $\bar{\mu}$ and $\bar{q}$, and $\bar{t}\in\mathbb{N}$ such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}\|q_{k}\|\leq\bar{q}$ via the bound in Lemma 3, ZO-HB has the same exponential decay rate as HB and converges to $\mathcal{B}_{r}(z^{e})$, and there exists a class-$\mathcal{K}$ function $\alpha$ such that
@@
Consider (22) and (23) with $P=TA$. Let $f$ be $\beta$-strongly convex with $L_{1}$-Lipschitz gradients, let the hypotheses of Lemma 4 be satisfied, and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of NAG. Then there exist $\bar{\mu},\bar{q}>0$ and $\bar{t}\in\mathbb{N}$ such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}E\|q_{k}\|\leq\bar{q}$ via the bound in Lemma 4, ZO-NAG has the same decay rate as NAG and converges to $\mathcal{B}_{r}(z^{e})$, and there exists a class-$\mathcal{K}$ function $\alpha$ such that
@@ 3.3 $L_{2}$ regularisation
In the preceding sections, we showed that strong convexity is sufficient to guarantee that the averaged ZO dynamics can be written as a bounded perturbation of the FO dynamics. When the objective $f$ is not strongly convex, one can introduce an $L_{2}$ regularisation term to induce the necessary contraction to recover the same perturbation framework, even when $f$ itself is not strongly convex.
@@
Consider (32), (34), and (35). Let $f$ be a possibly non-strongly-convex function with Lipschitz gradients with constant $L_{1}>0$, and let $h\in\bigl(0,\frac{\sqrt{2(1+c^{2})}-2c}{2L_{1}}\bigr)$. Then
$\sup_{k}\|q_{k}\|\leq h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-2L_{1}^{2}h^{2})}}.$ (36)
@@
Consider (32) and (34). Let $f$ be a possibly non-strongly-convex function with Lipschitz gradients with constant $L_{1}>0$, and let $h\in\bigl(0,\frac{\sqrt{2(1+c^{2})}-2c}{2L_{1}}\bigr),$ and for $r>0$ let $\mathcal{B}_{r}(z^{e})$ denote a convergence neighbourhood around the fixed point $z^{e}$ of GD. Suppose that GD applied to the regularised problem (31) converges to a fixed point $z^{e}$. Then there exist constants $\bar{\mu},\bar{q}>0$ and $\bar{t}\in\mathbb{N}$ (depending on $h$, $c$, $L_{1}$, and the Lyapunov function of the GD dynamics) such that, for any $\mu\in(0,\bar{\mu})$ and $t\geq\bar{t}$ satisfying $\sup_{k}\|q_{k}\|\leq\bar{q}$ via the bound in Lemma 5, ZO-GD has the same decay rate as GD and converges to $\mathcal{B}_{r}(z^{e})$ and there exists a class-$\mathcal{K}$ function $\alpha$ such that
$r=\alpha\left(h\frac{\mu}{2}L_{1}(n+3)^{3/2}+hL_{1}\sqrt{\frac{2h^{2}\sigma^{2}}{t(1-c^{2}-4cL_{1}h-2L_{1}^{2}h^{2})}}\right).$ (37)
The admissible parameter range in Lemma 5 implies that the regularised objective is strongly convex (Remark 8). Nevertheless, the framework provides a systematic way to apply ISS to originally non-convex objectives with explicit control over the trade-off between proximity to the original minimiser and neighbourhood size. Next, we illustrate the theoretical findings via numerical examples.